October 19th
------------

The retroreflecting optics has been installed by the LBTO crew in the morning around 9:00.

1) AO system performance
Two new reconstructor at BIN1 using 500 modes at 3lmabda/D and 2lambda/D has been measured. The two reconstructor has been tested succesfully achieving about 75% of SR. This results is in agreement with the expected performance of 0.74. The reconstructor at BIN2 has been checked and found consistent with the performance stated in the acceptance test table. In particular the BIN2 rec achieved a SR of 0.53 and 0.50 for mag 10.5 and 11.5. The expected performance are 0.58 and 0.56. The measured discrepancies could be due to telescope vibration. Some data reduction of BIN2 results is ongoing.

2) AO system SW

- The test of an automatic procedure to adjust the system overal gain has been started. The development of this procedure is ongoing.
- Some dubugging of the AdSec arbitrator has been done.

The retroreflector has been unmounted by the LBTO crew at 8:00PM. This has been done to try to close the actiity of the RFBG focal station commissioning. Howver due to incoming bad weather no observation has been possible. Finally during the night, between 1AM and 8AM, W. Seifert and N. Ageorges did some test to adjust the LUCIFER pupil position. The test was succesfull and they will let us know when they need additional telescope time.
Topic revision: r1 - 29 Oct 2010, AlfioPuglisi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback